Circular No. 6966

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980703 A. Levine, E. Morgan, and M. Muno, Massachusetts Institute of Tech-

nology (MIT), report on behalf of the RXTE ASM team at MIT and Goddard Space Flight Center: “A γ-ray burst that occurred on ∼ July 3.182 UT (BATSE trigger 6891) was in the field-of-view of two ASM cameras. The resulting detection yielded a position of α = 23^{h}59^{m}05^{s}, δ = +8^{o}33′.6 (equinox 2000.0), with an uncertainty of ∼ 4′ radius (90-percent confidence). The burst peak intensity was roughly 1.7 Crab (2–12 keV).”

SGR 1627–41 K. Hurley, Space Sciences Laboratory; T. Cline and P. Butterworth,

Goddard Space Flight Center, NASA; and E. Mazets and S. Golenetskii, Ioffe Institute, on behalf of the Ulysses and Konus-Wind teams, report: “We have triangulated the position of this soft γ-ray repeater using seven Konus-Ulysses events between June 17 and 22. A typical annulus has a center at α = 22^{h}03^{m}04^{s}, δ = −9^{o}49′.8 (equinox 2000.0), and a radius of 77^{o}.247 ± 0^{o}.013. Around the position of the supernova remnant (SNR) G337.0–0.1, this annulus nests within the one reported previously (IAUC 6948), but is considerably narrower. Around this position, the seven annuli give results that are mutually consistent to better than 10′′. The positions of the annuli suggest that, if the source of the bursts is indeed within the SNR, it lies between the two radio lobes. An image may be found at

http://ssl.berkeley.edu/ipn3/sgr1627-41/.”

JUNE BOOTID METEORS 1998 P. Brown and W. K. Hocking, University of Western Ontario, write in

regard to the meteor activity reported on IAUC 6954: “Radar observations with the Skiymet meteor radar (35.24 MHz) in Saskatoon, SK, show a strong increase in radar rates centered on June 27.60 ± 0.04 UT. Using radiant mapping techniques (Jones and Morton 1982, MNRAS 200, 281) in conjunction with a 5-element interferometer, a radiant centered at α = 228^{o} ± 3^{o}, δ = +54^{o} ± 3^{o} (equinox 2000.0) is evident from meteor echo data, though the radiant area is quite diffuse (with an additional radiant activity center near α = 219^{o}, δ = +61^{o}). The peak shower flux (assuming a mass index of 2.0) to the radar limiting absolute magnitude of +7.5 was 0.12 ± 0.02 meteoroids km^{−2} hr^{−1}. The corresponding orbit is similar to that of comet 7P/Pons-Winnecke and suggests a recurrence of the periodic June Bootid shower associated with that comet.”

1998 July 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green